# Modular Recurrence in Contextual MDPs for Universal Morphology Control

### v1

@@ 2.2 Partially Observable Context
In this paper, we recognize and exploit the fact that relevant contextual information can be partially observable. Robot observations received by the agent contain certain contextual properties of the robot morphology and topology, for example the mass and shape of the different limbs, and their initial positions. The features that are provided, however, depend on what information is available and whether they can be effectively processed in the agent architecture. A lot of (possibly) relevant contextual information is not available, such as the friction and damping of limbs. Additionally, the true adjancency matrix that describes the organization of the limbs can not be provided effectively to an agent with a modular architecture, as pointed out by Xiong et al. (2023). Therefore, topological information is provided rather implicitly instead, through the position of a limb with respect to its parent limb. Lastly, we do not have an adequate way of providing the agent with more abstract features, such as the influence of different limbs on each other during movement. It is only through interaction with the environment that the agent can infer such features to perform optimally in the current task. It stands to reason that state-of-the-art approaches, which rely on provided context features for generalization, can only learn these interactions by overfitting to the training contexts.
@@ 5.3 Zero-Shot Generalization to Different Dynamics and Kinematics
Training Flat Terrain $3697\pm 137$ $3819\pm 61$ $\mathbf{4230\pm 101}$ $\mathbf{4186\pm 58}$
@@
Testing Flat Terrain $1384\pm 92$ $1678\pm 87$ $1455\pm 94$ $\mathbf{1829\pm 97}$
@@
5.4 Zero-Shot Generalization to Unseen Robot Topologies
After training on the 100 training robots, the methods were evaluated on the 70 test robots with unseen topologies. The averaged returns of the different methods in the four environments are denoted in Table 1. In each of the environments, ModuMorph and/or R-MoMo dominate training, but R-MoMo significantly outperforms the other methods in zero-shot generalization to the test robots. Interestingly, in all environments (except Incline) ModuMorp and R-MoMo perform roughly the same during training, but the recurrent version performs significantly better in zero-shot generalization. This indicates that R-MoMo does not just learn a better policy, but a policy that generalizes much better to unseen robots. The recurrent version of MetaMorph shows a similar improvement on test robots for all environments, but also outperforms its baseline in more training environments, which makes it less clear that the advantage comes from better generalization. In comparison, R-MeMo performs roughly on a par with vanilla ModuMorph on test robots.
Returns in each environment can range from very low ($<0$) to very high ($>4000$) values. Averaged returns over 70 test robots could therefore be misleading, as differences can be caused by only a small number of test robots. We therefore additionally report the per-robot difference in return (averaged over 5 seeds) between R-MoMo and ModuMorph on the Flat Terrain environment in Figure 6. These results show a consistent improvement over a majority of the test robots, illustrating the increased generalizability. Per-robot performance comparisons for the other environments can be found in Appendix B.
@@ 6 Discussion
This work explores improved generalization in multi-robot control. It was hypothesized that modular recurrence could allow the agent to infer relevant unobservable context to improve performance. The results have shown a consistent increase in generalization performance when such memory-mechanism was introduced across different environments, for robots with different dynamics, kinematics, and topologies. Whether the agent actually uses the RNN to infer relevant contextual features remains to be investigated. A limitation of the explored recurrent architecture, is the fact that hidden states have to be stored for each limb. Scaling to robots with a large number of limbs requires therefore more memory. Nonetheless, the combination of modular recurrence with transformers can be promising in domains with a graph-like structure of which the relevant properties are only partially observable.

### v2

@@ 2.2 Partially Observable Context
In this paper, we recognize and exploit the fact that relevant contextual information can be partially observable. Robot observations received by the agent contain certain contextual properties of the robot morphology and topology, for example the mass and shape of the different limbs, and their initial positions. The features that are provided, however, depend on what information is available and whether they can be effectively processed in the agent architecture. A lot of (possibly) relevant contextual information is not available, such as the friction and damping of limbs. Additionally, the true adjacency matrix that describes the organization of the limbs can not be provided effectively to an agent with a modular architecture, as pointed out by Xiong et al. (2023). Therefore, topological information is provided rather implicitly instead, through the position of a limb with respect to its parent limb. Lastly, we do not have an adequate way of providing the agent with more abstract features, such as the influence of different limbs on each other during movement. It is only through interaction with the environment that the agent can infer such features to perform optimally in the current task. It stands to reason that state-of-the-art approaches, which rely on provided context features for generalization, can only learn these interactions by overfitting to the training contexts.
@@ 3.1 Universal Morphology Control
Steps towards robotic applications in the real world also exploit such modular transformer-based architectures for generalizable and transferrable control. Several recent works incorporate additional information about the robot topology through a graph encoding or sparse attention matrices (Patel and Song, 2024; Sferrazza et al., 2024) for improved generalization in simulation and the real world. Fine-tuning and policy distillation approaches for effective generalization have also been proposed for generalization (Przystupa et al., 2025; Xiong et al., 2024). Lastly, more complex modular architectures were introduced for effective transfer to unseen robots in simulation and the real world (Bohlinger et al., 2024; Li et al., 2024), although evaluating on smaller sets of, and relatively similar robot morphologies. None of these approaches consider this generalization problem as partially observable, where information about the robot to be controlled can be inferred.
@@ 5.3 Zero-Shot Generalization to Different Dynamics and Kinematics
Training
Flat Terrain $3697\pm 137$ $3819\pm 61$ $\mathbf{4230\pm 101}$ $\mathbf{4186\pm 58}$
@@
Testing
Flat Terrain $1384\pm 92$ $1678\pm 87$ $1455\pm 94$ $\mathbf{1829\pm 97}$
@@
5.4 Zero-Shot Generalization to Unseen Robot Topologies
After training on the 100 training robots, the methods were evaluated on the 70 test robots with unseen topologies. The averaged returns of the different methods in the four environments are denoted in Table 1. In each of the environments, ModuMorph and/or R-MoMo dominate training, but R-MoMo significantly outperforms the other methods in zero-shot generalization to the test robots. Interestingly, in all environments (except Incline) ModuMorph and R-MoMo perform roughly the same during training, but the recurrent version performs significantly better in zero-shot generalization. This indicates that R-MoMo does not just learn a better policy, but a policy that generalizes much better to unseen robots. The recurrent version of MetaMorph shows a similar improvement on test robots for all environments, but also outperforms its baseline in more training environments, which makes it less clear that the advantage comes from better generalization. In comparison, R-MeMo performs roughly on a par with vanilla ModuMorph on test robots.
@@ 5.4 Zero-Shot Generalization to Unseen Robot Topologies
Returns in each environment can range from very low (<0<0) to very high (>4000>4000) values. Averaged returns over 70 test robots could therefore be misleading, as differences can be caused by only a small number of test robots. We therefore additionally report the per-robot difference in return (averaged over 5 seeds) between R-MoMo and ModuMorph on the Flat Terrain environment in Figure 6. These results show a consistent improvement over a majority of the test robots, illustrating the increased generalizability. Per-robot performance comparisons for the other environments can be found in Appendix B.
@@ 6 Discussion
This work explores improved generalization in multi-robot control. It was hypothesized that modular recurrence could allow the agent to infer relevant unobservable context to improve performance. The results have shown a consistent increase in generalization performance when such memory-mechanism was introduced across different environments, for robots with different dynamics, kinematics, and topologies. Whether the agent actually uses the RNN to infer relevant contextual features remains to be investigated. A limitation of the explored recurrent architecture, is the fact that hidden states have to be stored for each limb. Scaling to robots with a large number of limbs requires therefore more memory. Besides, the sequential processing of RNNs results in longer training times (although minimized by efficient batch processing through chunking episodes). An interesting direction for future research would therefore be to investigate a more efficient memory-mechanism in the architecture. Nonetheless, the combination of modular recurrence with transformers can be promising in domains with a graph-like structure of which the relevant properties are only partially observable.
Acknowledgements
This project was funded by the Delft AI initiative within the BIOlab. This work was also partially funded by the Dutch Research Council (NWO) project Reliable Out-of-Distribution Generalization in Deep Reinforcement Learning with project number OCENW.M.21.234. DB acknowledges support from an ERC Starting Grant (850818 - MULTI-VIsion) and the Convergence Health and Technology (Integrative Neuromedicine flagship).